Baab Corporation is a manufacturing firm that uses job-order costing. The companys inventory balances were as follows at theRequired A Required B Required Prepare a schedule of cost of goods manufactured. (Round your intermediate calculations to 2 dRequired A Required B Required C Was the overhead underapplied or overapplied? By how much?Prepare an income statement for the year. The company closes any underapplied or overapplied overhead to Cost of Goods Sold.
Answer #1

Part.1

Baab Corporation
Schedule of Cost of Goods Manufactured
Direct Material
Opening 14350
Add Purchased 315350
Total Raw Material Available 329700
Less: Closing Raw Material -22350
Less: Indirect Raw Material used in Production -26700
Raw Material used in production (A) 280650
Direct Labour       (B) 377350
Factory Overhead applied ($7.7[Note.1]*34070) (C ) 262339
Total manufacturing costs (A+B+C) 920339
Add Opening Work in Progress 27350
Less: Closing Work in Progress -9350
Cost of goods manufactured 938339
Note.1 Factory Overhead Recovery Rate = Budgeted Factory Overhead/Machine Hours
=256795/33350 = $       7.70

Part. 2

Computation of Manufacturing Overhead Incurred
Indirect Material 26700
Indirect Labour 96350
Factory Utilities Cost 10350
Depreciation 131000
Total Manufacturing Overhead 264400
Manufacturing Overhead Under or Over applied = Actual Manufacturing Overhead Incurred - Manufacturing Overhead applied
= 264400 - 262339 = 2061 Underapplied

Part 3

Baab Corporation
Income Statement
Sales 1267000
Add: Closing Finised Goods 77350
Less: Opening Finished Goods -62350
15000
Less: Selling & Administrative Expense
Administrative Salaries 172350
Depreciation relating to the selling, general & administrative activities 17000
Selling Costs 147350
-336700
Less: Underapplied Overheads -2061
Less: Cost of Goods Manufactured -938339
Profit/Loss 4900